OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

ganeshie8 (ganeshie8):

use dot product u.v = ||u|| ||v|| cos\(\theta\)

ganeshie8 (ganeshie8):

<-5, -4> . <-4, -3> = ||<-5, -4> || ||<-4, -3> || cos \(\theta\) solve \(\theta\)

OpenStudy (anonymous):

I keep getting 5 squareroot 41

ganeshie8 (ganeshie8):

yes i also got 5 squareroot41 :) <-5, -4> . <-4, -3> = ||<-5, -4> || ||<-4, -3> || cos \(\theta\) 20+12 = \(\sqrt{41} \times 5\) cos \(\theta\) 32 = \(5\sqrt{41} \) cos \(\theta\)

ganeshie8 (ganeshie8):

you still need to solve \(\theta\)

ganeshie8 (ganeshie8):

\(\large \frac{32}{5\sqrt{41}} = \cos \theta\)

ganeshie8 (ganeshie8):

use ur calculator and find \(\theta\)

ganeshie8 (ganeshie8):

\(0.999512 = \cos \theta\) \(\cos \theta = 0.999512 \) can u find \(\theta\) ?

OpenStudy (anonymous):

Wait, 0.9 is the answer, if I am not mistaken.

ganeshie8 (ganeshie8):

nopes

ganeshie8 (ganeshie8):

when u have, \(\cos \theta = 0.999512\) you must take cos inverse both sides, to isolate \(\theta\)

ganeshie8 (ganeshie8):

\(\cos \theta = 0.999512\) \(\cos^{-1}\cos \theta = \cos^{-1}0.999512\) \( \theta = \cos^{-1}0.999512\)

ganeshie8 (ganeshie8):

do you knw how to take \(\cos^{-1} 0.999512\) in your calcualtor ? :)

ganeshie8 (ganeshie8):

you may use wolframalpha :- http://www.wolframalpha.com/input/?i=arccos%280.99951207%29

OpenStudy (anonymous):

1.79 = 1.8.

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

Once again, Thank you, very much.

ganeshie8 (ganeshie8):

you're wlcme :)